OpenStudy (daniellelovee):

Farmer Brown had 6 water wells drilled on his land. The drillers charge by the foot in terms of the depth of the well, plus fees for labor, parts, and fuel. The points on the scatter plot below show the cost of the six wells as well as their depth. The red line has been fit to the data for estimation and prediction. Which of the following statements are valid representations of the meaning of the 1000 in the equation of the line?

OpenStudy (daniellelovee):

Valid Not Valid A 1000 is the predicted cost of a well without any charge based on the depth of the well. B 1000 is the slope of the prediction line. C 1000 is the y-intercept of the prediction line. D 1000 is the predicted number of dollars the cost increases by for every additional foot of depth.

OpenStudy (daniellelovee):

so A is false

OpenStudy (daniellelovee):

B is false C is true and D is true?

Directrix (directrix):

On A, put 0 in for x in y = 5x + 1000 and see what you get.

OpenStudy (daniellelovee):

i get 1000

OpenStudy (daniellelovee):

@Directrix internet problems again?

Directrix (directrix):

No. I asked you a question and saw no reply. I went off to help an OP.

Directrix (directrix):

If you get 1000, then why is question A false?

OpenStudy (daniellelovee):

because for the last one I got I got 1400 and you said it was false and it was correct

Directrix (directrix):

I need to look at that problem. Which one of the golf course problems was it? A, B, C, or D?

OpenStudy (daniellelovee):

yes the one that talked about the puff and the green

Directrix (directrix):

Which question? A?

OpenStudy (daniellelovee):

yes

Directrix (directrix):

It's your work so go with what you think. For this, I have: T F T F

OpenStudy (daniellelovee):

ok thank you :)

Directrix (directrix):

On A, put 0 in for x in y = 5x + 1000 to get 1000. Startup cost is 1000 before the drilling starts.

OpenStudy (daniellelovee):

I'm just missing 2 more from this lesson to have 15 questions and have 90%+

Directrix (directrix):

That is why I chose T for part A of this question.

OpenStudy (daniellelovee):

ohh ok that makes sense so is like a regular line and y=mx+b equation?

Directrix (directrix):

Did you miss the entire golf course problem because one of the 4 associated questions was wrong?

Directrix (directrix):

Yes, and I did the same sort of reasoning on part A of the golf course question. So, I don't know what to tell you.

OpenStudy (daniellelovee):

i don't know how I didn't get that either I'm sorry :(

OpenStudy (daniellelovee):

and your answers were correct for the golf question

Directrix (directrix):

y = 5x + 1000 y = 5*0 + 1000 y = 1000 The tricky thing here is how the question is worded: A 1000 is the predicted cost of a well without any charge based on the depth of the well. 1000 is NOT the predicted cost of the entire well project. Therefore, a person could argue False on part A.

OpenStudy (daniellelovee):

it showed as correct but that website has some issues most of the time even when they give you the step by step solution they tell you something like 1. 1+2X=7 and then do 7-1=? and then when you put 6 they will tell you wrong and then on the next step it would show as 2.7-1=6 6/2=?

OpenStudy (daniellelovee):

so the have to still fix stuff but your answers have been correct so far thank you very much :)

Directrix (directrix):

The A part question on the well problem is not analogous to the A part question on the golf course problem. This is a weird-i-o question. We'll just have to move on and do our best.

OpenStudy (daniellelovee):

yeah I agree :)

Directrix (directrix):

Did you check the well digging answers?

OpenStudy (daniellelovee):

yes :)

Directrix (directrix):

Alrighty, then. Time to move on.

OpenStudy (daniellelovee):

yup :)